OpenStudy (anonymous):

Check my answer, please?: Find the volume of a sphere with surface area equal to 100π ft2.

OpenStudy (anonymous):

A=π⅓(6V)⅔=π⅓·(6·36)⅔≈52.72531

OpenStudy (anonymous):

@amistre64

OpenStudy (amistre64):

\[V=4\pi r^3\] \[S_a=\frac43\pi~r^2\] these seem familiar

OpenStudy (amistre64):

find r from Sa to determine V

OpenStudy (amistre64):

sqrt(S*3/(4pi)) = r V = 4pi * S*3/(4pi) * sqrt(S*3/(4pi)) V = S*3* sqrt(S*3/(4pi)) V = 300pi* sqrt(100*3/4) V = 300pi* 10/2 * sqrt(3) V = 1500pi sqrt(3) maybe?

OpenStudy (amistre64):

i got those backwards :/

OpenStudy (anonymous):

Oh okay. From the top! c:

OpenStudy (amistre64):

S = 4pi r^2 ... V = 4pi r^2/3

OpenStudy (amistre64):

** r^3/3

OpenStudy (amistre64):

so, it looks like V = S * r/3

OpenStudy (amistre64):

r = sqrt(S/4pi) = 5 in this case?

OpenStudy (anonymous):

So √(100/4π)?

OpenStudy (amistre64):

S = 100 pi sqrt(100pi/4pi) = sqrt(100/4) = 10/2 = 5 = r

OpenStudy (amistre64):

V = S*r/3 = 500pi/3 seems about more righter lol

OpenStudy (anonymous):

Okay so given the radius now, we would do 4/3π(5)^2 ?

OpenStudy (amistre64):

5^3 but yes

OpenStudy (amistre64):

or the simplifed S*r/3

OpenStudy (anonymous):

100π*5/3?

OpenStudy (anonymous):

@ParthKohli

OpenStudy (anonymous):

@mathstudent55

OpenStudy (anonymous):

using the above formulas i get r = sqrt(75) ft V =2598pi is that what you get?